# History, Development, and Principles of Large Language Models—An Introductory Survey

### v2

@@ Abstract
Language models serve as a cornerstone in natural language processing (NLP), utilizing mathematical methods to generalize language laws and knowledge for prediction and generation. Over extensive research spanning decades, language modeling has progressed from initial statistical language models (SLMs) to the contemporary landscape of large language models (LLMs). Notably, the swift evolution of LLMs has reached the ability to process, understand, and generate human-level text. Nevertheless, despite the significant advantages that LLMs offer in improving both work and personal lives, the limited understanding among general practitioners about the background and principles of these models hampers their full potential. Notably, most LLMs reviews focus on specific aspects and utilize specialized language, posing a challenge for practitioners lacking relevant background knowledge. In light of this, this survey aims to present a comprehensible overview of LLMs to assist a broader audience. It strives to facilitate a comprehensive understanding by exploring the historical background of language models and tracing their evolution over time. The survey further investigates the factors influencing the development of LLMs, emphasizing key contributions. Additionally, it concentrates on elucidating the underlying principles of LLMs, equipping audiences with essential theoretical knowledge. The survey also highlights the limitations of existing work and points out promising future directions.
@@ 1. INTRODUCTION
Language stands as humanity’s most potent tool, enabling the expression of thoughts and emotions and facilitating communication with others [pinker2003language, turing2009computing]. However, machines lack the intrinsic ability to comprehend and communicate in human language, necessitating powerful Artificial Intelligence (AI) algorithms to acquire such capabilities. The ultimate aim of AI research is the creation of machines capable of reading, writing, and conversing like humans [dwivedi2023so, zhao2023survey, jin2023rethinking]. Natural language processing (NLP) emerges as the AI branch dedicated to realizing this goal. At the core of NLP are language models, tasked with predicting or generating the probability or likelihood of linguistic units (e.g., words, phrases, sentences) based on context. The evolutionary stages of language models unfold chronologically from the initial statistical language models (SLMs) to subsequent neural language models (NLMs), progressing to pre-trained language models (PLMs), and reaching the current state of large language models (LLMs). Specifically, SLMs employ simple probability distributions to model word sequences, while NLMs utilize neural networks to grasp complex patterns and representations of language. PLMs leverage large-scale corpora and self-supervised learning to capture general linguistic knowledge, and LLMs extend PLMs by incorporating massive data, computation, and algorithms, resulting in more expressive, sophisticated, and adaptable language models [fu2022does, wei2022emergent].
@@
In recent years, significant advancements in NLP have been achieved with LLMs, such as the generative pre-trained transformer (GPT) series of models [radford2018improving, radford2019language, brown2020language, achiam2023gpt]. Extensively trained on textual data, these models demonstrate the ability to generate human-level texts and perform language-based tasks with exceptional precision. Despite these advancements having greatly benefited various aspects of work and life, LLMs might not be widely understood by practitioners, particularly those without a background in NLP. To provide practitioners with a fundamental understanding of LLMs, this review introduces it across six dimensions: history, development, principles, applications, drawbacks and future directions. This survey distinguishes itself from other reviews on PLMs or LLMs [zhao2023survey, liu2023pre, han2021pre, shanahan2024talking], which often concentrate on specific aspects using technical language; this survey aims to elucidate LLMs and their principles in a more accessible manner, intending to maximize its full potential.
@@
The subsequent sections of this review are structured as follows: Section 2 details the history of LLMs and the contributing factors behind their rapid growth, covering the evolution of language models, increased data diversity, computational advancements, and algorithmic innovations. Section 3 offers an overview of LLMs principles, using the GPT family of models as an accessible example for better comprehension. Moving on, Section 4 delves into the diverse applications of LLMs, spanning drug discovery, finance, medical, and legal domains. Section 5 critically examines the drawbacks of current state-of-the-art LLMs, and finally, Section 6 concludes the survey.
@@ 2. HISTORY AND DEVELOPMENT OF LARGE LANGUAGE MODELS
In this section, we will explore the history of the language model and analyze the developmental dynamics and influences shaping LLMs.
@@ 2.1 History of LLMs
To better understand LLMs, this section will follow the developmental stages of language models and introduce SLMs, NLMs, PLMs, and LLMs. Figure 1 provides a visual map of the history of the development of the language model.
@@
  • Statistical Language Model: SLMs [jelinek1998statistical, stolcke2002srilm, rosenfeld2000two] originated in the 1990s as mathematical models addressing contextually relevant properties of natural language from a probabilistic statistical perspective. The essence of statistical language modeling lies in ascertaining the probability of a sentence occurring within a text. Considering $S$ as the sentence “I am very happy”, $P(\omega_{i})$ signifies the probability of the $i$-th word in the sentence: $\omega_{1}$ as “I”, $\omega_{2}$ as “am”, $\omega_{3}$ as “very”, and $\omega_{4}$ as “happy”. Now, the objective is to ascertain the likelihood of $S$ appearing in the text, denoted as $P(S)$:
@@
  • Neural Language Models: NLMs [bengio2000neural, mikolov2010recurrent, kombrink2011recurrent] leverage neural networks to predict the probabilities of subsequent words within sequences. They effectively handle longer sequences and mitigate the limitations associated with small $n$ in SLMs. Before delving into neural networks, let’s grasp the concept of word vectors. Humans effortlessly comprehend word meanings. For instance, “cat” and “dog” share closer semantic connections than “cat” and “tree” since they both represent animals. But how does a computer accomplish this? Computers operate using binary code — 0s and 1s — so human language needs translation into binary. Word vectors can accomplish this, which are numerical representations of human language, where each word corresponds to a distinct vector. These vectors usually possess fixed dimensions and can simulate word relationships through the angles between them. An intriguing example is the angle between the word vectors for “cat” and “dog” which is smaller than the angle between the word vector for “cat” and “tree”. Another illustrative example is “$M\ essi-Argentina+P\ ortugal=Cristiano\ Ronaldo$”. Word2Vec [mikolov2013distributed] is a widely recognized tool for computing word vectors. Its function involves converting words into dense numerical representations, making words with similar semantics closer together in the vector space.
@@
    The efficacy of Word2Vec is notably influenced by its utilization of neural networks, which mirrors the structure and functionality of the human brain, comprising interconnected simple units known as neurons organized into different layers. Figure 2 provides a visualization of a simple NLM structure, composed of input, hidden, and output layers. Within this structure, each word vector $x_{t}\in\mathbb{R}^{m}$, the matrix $W_{xh}\in\mathbb{R}^{m(n-1)\times h}$, and the matrix $W_{ho}\in\mathbb{R}^{h\times|V|}$, where $h$ signifies the number of neurons in the hidden layer, and $V$ is the vocabulary containing all words recognized by the model. NLMs operate akin to the n-gram concept, assuming that the probability of each word depends only on its previous n-1 words. The first layer—the input layer—concatenates the word vectors of n-1 words, forming a vector $X\in\mathbb{R}^{m\times(n-1)}$. Subsequently, the second layer—the hidden layer—derives the output $H$ by applying a non-linear activation function, like sigmoid or tanh, to the matrix product of $X$ and $W_{xh}$. Following this, the third layer—the output layer—aims to forecast the subsequent word based on the hidden layer’s output. With $|V|$ neurons within the output layer, the resulting output vector $O\in\mathbb{R}^{|V|}$ is computed by multiplying $H$ and $W_{ho}$. This resultant vector then undergoes the Softmax function, producing a vector $O’$ ’ containing the probability value assigned to each word. The Softmax function, concerning the output of the $i$-th neuron, is defined as follows:
    $O_{i}^{\prime}=\operatorname{Softmax}\left(O’_{i}\right)=\frac{\exp\left(O_{i}\right)}{\sum_{j=1}^{|V|}\exp\left(O_{j}\right)},$ (4)
@@
    where $O_{i}^{\prime}$ denotes the output value of the $i$-th node, and $|V|$ represents the count of output nodes, corresponding to the classification categories. Utilizing the Softmax function allows the transformation of output values from multiclassification into a probability distribution, ensuring a sum of 1 within the range [0, 1].
@@
  • Pre-trained Language Model: PLMs undergo initial training using an extensive volume of unlabeled text, enabling them to grasp fundamental language structures such as vocabulary, syntax, semantics, and logic — a phase termed pre-training. Subsequently, this comprehensive language model can be applied to various NLP tasks like machine translation, text summarization, and question-answering systems. To optimize its performance, models need to be trained a second time on a smaller dataset customized for a specific downstream task — a phase known as fine-tuning. This is the “pre-training and fine-tuning” learning paradigm. We can use a visual example to understand the “pre-training and fine-tuning”, as follows: in martial arts novels, a person who wants to become a martial arts master needs to have a solid foundation of internal martial arts, which can be acquired by training on a large variety of techniques. Then, the person can learn a specific skill, such as a sword or a palm strike, and master it quickly and effectively by applying internal martial arts principles. A large number of studies on PLMs have been built on this paradigm, which introduces different architectures [devlin2018bert, fedus2022switch] (e.g., GPT-2 [radford2019language] and Bert [devlin2018bert]).
@@
  • Large Language Model: LLMs are trained on massive text corpora with tens of billions (or more) of parameters, such as GPT-3 [brown2020language], GPT-4 [achiam2023gpt], PaLM [chowdhery2023palm], and LLaMA [touvron2023llama]. The goal of LLMs is to enable machines to understand human commands and adherence to human values. Their hallmark lies in the consolidation of two stages: initial pre-training on a vast general-purpose corpus followed by alignment with human values, rather than transitioning to a different domain. LLMs exhibit remarkable adaptability compared to PLMs, transitioning from specialized to general-purpose models. The substantial increase in model size, dataset volume, and computational prowess has resulted in significant enhancements across various tasks and unveiled remarkable capabilities absent in smaller models. For example, GPT-3 has the capability to leverage contextual information, a functionality that GPT-2 [radford2019language] lacks. This means that when GPT-3 is prompted with task-related examples, it utilizes them to enhance its problem-solving capabilities. The number of parameters for LLMs typically exceeds a hundred billion, and the training data is usually in the range of a few hundred GB to a few TB. A concrete example is that there are multiple versions of the GPT-2 model, with the largest version having 1.5 billion parameters and using 40GB of text data for training. In contrast, the largest version of the GPT-3 model has 175 billion parameters and uses 570GB of text data for training. This example illustrates the significant discrepancy between LLMs and PLMs concerning parameter count and training data volume.
@@ 2.2 Factors Propelling Large Language Models
  • Data Diversity: data diversity has emerged as a crucial catalyst for the development of LLMs. Recent years have witnessed a surge in the availability and diversity of extensive internet-based data sources, furnishing these models with training data containing rich linguistic and worldly insights. The quality and origins of this data significantly influence the efficacy and capacity of LLMs. Earlier studies trained language models on a single domain of text, such as news articles [jozefowicz2016exploring], Wikipedia [merity2016pointer], or fiction books [kiros2015skip]. However, the proliferation of vast and heterogeneous text corpora across the web—including blogs, social media, forums, reviews, among others—has spurred researchers to pivot towards training models on multi-domain texts. This shift enhances the model’s aptitude for generalization and multitasking. An exemplary instance is Meta AI’s open-source large language model, LLaMA [touvron2023llama], trained on publicly accessible datasets like CommonCrawl [^1] https://commoncrawl.org/get-started, C4 [raffel2020exploring], Github [^2] https://github.com/, Wikipedia [^3] https://huggingface.co/datasets/wikipedia, Books [gao2020pile], ArXiv [^4] https://arxiv.org/, and StackExchange [^5] https://stackexchange.com/.
@@
  • Computational Advancement: LLMs are neural network models with a large number of parameters. With the progression of deep learning (DL) technology in recent years, the scale and effectiveness of LLMs have notably escalated. Simultaneously, they present serious challenges due to their substantial computational demands [kaplan2020scaling]. Computational power signifies a computer system’s capability to execute computational tasks, often quantified in terms of floating-point operations per second. According to OpenAI calculations [^6] https://openai.com/research/ai-and-compute, since 2012, the amount of computation used for global AI training has shown exponential growth, doubling every 3.43 months on average, and now the amount of computation has expanded 300,000 times, far outpacing the rate of computational power growth. Fortunately, hardware innovators like Nvidia continually invent specialized devices such as GPUs, TPUs, and NPUs with amplified computational prowess, LLMs can be trained quickly. For instance, the launch of GPT-3 by OpenAI in May 2020 [brown2020language] underscores this reliance on robust hardware. Training GPT-3 on a single NVIDIA Tesla V100 GPU would require an estimated 355 years. Conversely, leveraging 1024×A100 GPUs, researchers estimated the potential to train GPT-3 in as little as 34 days. This exemplifies the indispensability of potent hardware in facilitating the development and training of LLMs.
@@
  • Algorithmic Innovation: algorithmic innovation stands as the pivotal driving force behind LLMs, shaping their structure and functionality. From the initial rule- and statistics-based approach, to the later DL-based approach, the algorithms of language models have continuously evolved and improved. Presently, all LLMs are based on the transformer architecture [vaswani2017attention], which employs a self-attention mechanism (see discussion in Section 3). This architecture offers distinct advantages over traditional recurrent [mikolov2010recurrent] and convolutional neural networks [lecun1998gradient], enabling fully parallel computation and adeptly capturing long-distance dependencies. Moreover, there exist numerous iterations and enhancements of the transformer, such as Transformer-XL [cui2023chatlaw], XLNet [yang2019xlnet], ALBERT [lan2019albert], each targeting specific aspects for optimization. These variations aim to enhance the attention mechanism, refine pre-training objectives, and curtail parameter counts, thereby advancing the transformer architecture in diverse dimensions.
3. PRINCIPLES OF LARGE LANGUAGE MODELS

### v3

@@ Abstract
Language models serve as a cornerstone in natural language processing (NLP), utilizing mathematical methods to generalize language laws and knowledge for prediction and generation. Over extensive research spanning decades, language modeling has progressed from initial statistical language models (SLMs) to the contemporary landscape of large language models (LLMs). Notably, the swift evolution of LLMs has reached the ability to process, understand, and generate human-level text. Nevertheless, despite the significant advantages that LLMs offer in improving both work and personal lives, the limited understanding among general practitioners about the background and principles of these models hampers their full potential. Notably, most LLM reviews focus on specific aspects and utilize specialized language, posing a challenge for practitioners lacking relevant background knowledge. In light of this, this survey aims to present a comprehensible overview of LLMs to assist a broader audience. It strives to facilitate a comprehensive understanding by exploring the historical background of language models and tracing their evolution over time. The survey further investigates the factors influencing the development of LLMs, emphasizing key contributions. Additionally, it concentrates on elucidating the underlying principles of LLMs, equipping audiences with essential theoretical knowledge. The survey also highlights the limitations of existing work and points out promising future directions.
keywords:
Large Language Model, Natural Language Processing, Artificial Intelligence, Language Model
@@ 1. Introduction
Language stands as humanity’s most potent tool, enabling the expression of thoughts and emotions and facilitating communication with others [1, 2]. However, machines lack the intrinsic ability to comprehend and communicate in human language, necessitating powerful Artificial Intelligence (AI) algorithms to acquire such capabilities. The ultimate aim of AI research is the creation of machines capable of reading, writing, and conversing like humans [3, 4, 5]. Natural language processing (NLP) emerges as the AI branch dedicated to realizing this goal. At the core of NLP are language models (LMs), tasked with predicting or generating the probability or likelihood of linguistic units (e.g., words, phrases, sentences) based on context. The evolutionary stages of LMs unfold chronologically from the initial statistical language models (SLMs) to subsequent neural language models (NLMs), progressing to pre-trained language models (PLMs), and reaching the current state of large language models (LLMs). Specifically, SLMs employ simple probability distributions to model word sequences, while NLMs utilize neural networks to grasp complex patterns and representations of language. PLMs leverage large-scale corpora and self-supervised learning to capture general linguistic knowledge, and LLMs extend PLMs by incorporating massive data, computation, and algorithms, resulting in more expressive, sophisticated, and adaptable LMs [6, 7].
@@
In recent years, significant advancements in NLP have been achieved with LLMs, such as the generative pre-trained transformer (GPT) series of models [8, 9, 10, 11]. Extensively trained on textual data, these models demonstrate the ability to generate human-level texts and perform language-based tasks with exceptional precision. Despite these advancements having greatly benefited various aspects of work and life, LLMs might not be widely understood by practitioners, particularly those without a background in NLP. To provide practitioners with a fundamental understanding of LLMs, this review introduces it across six dimensions: history, development, principles, applications, drawbacks and future directions. This survey distinguishes itself from other reviews on PLMs or LLMs [4, 12, 13, 14], which often concentrate on specific aspects using technical language; this survey aims to elucidate LLMs and their principles in a more accessible manner, intending to maximize its full potential.
To achieve a comprehensive survey of LLMs, we employed a meticulous review process guided by a structured search strategy. We searched databases such as IEEE Xplore, ACM Digital Library, Scopus, and Google Scholar using keywords including “large language models”, “large language models survey”, and “large language models for X” where X includes software engineering, drug discovery, finance, medical, legal, and education. Our review focused on publications from major conferences and journals about NLP and LLMs’ applications, including EMNLP [15, 16, 17, 18], ACL [19, 20, 21, 22], ICLR [23, 24, 25, 26, 27, 28, 29, 30], ICML [31, 32, 33, 34], ICSE [35, 36], ICAIF [37, 38], ICDM [39, 40, 41], IJCAI [42, 43, 44], AAAI [45, 46, 47], ECAI [48, 49, 50], BEA [51], Financial Innovation [52], PLOS Digital Health [53, 54], Applied Sciences [55], Interspeech [56, 57, 58], Journal of Machine Learning Research [59, 60, 61, 62], Journal of AI [63], etc., with an emphasis on studies from the past decade. Additionally, we reviewed references from selected papers and cross-checked our findings against existing LLMs surveys [4, 64, 65, 66]. Our process involved identifying relevant studies and applying screening criteria based on LLMs relevance, citation count, and the quality of the publication venue. Papers with minimal contributions or lower relevance were excluded. The remaining studies were then evaluated for methodological rigor and clarity of conclusions. Key data from these studies were extracted and synthesized to highlight trends, gaps, and emerging themes in LLM research. This synthesis offers a comprehensive overview of current advancements, ongoing challenges, and future research directions, providing valuable insights for both academic and industrial applications.
@@
The subsequent sections of this review are structured as follows: Section 2 details the history of LLMs and the contributing factors behind their rapid growth, covering the evolution of LMs, increased data diversity, computational advancements, and algorithmic innovations. Section 3 provides an overview of the principles of LLMs, using the GPT family of models as an accessible example to enhance understanding of LLMs’ principles. Additionally, we present a comparison of state-of-the-art models across various types to assist readers in selecting the most suitable model for their specific applications. Moving on, Section 4 explores the wide-ranging applications of LLMs across various industries and professional domains. Section 5 critically examines the drawbacks of current state-of-the-art LLMs, and finally, Section 6 concludes the survey.
@@ 2. HISTORY AND DEVELOPMENT OF LARGE LANGUAGE MODELS
In this section, we will explore the history of the LMs and analyze the developmental dynamics and influences shaping LLMs.
@@ 2.1 History of LLMs
To better understand LLMs, this section will follow the developmental stages of LMs and introduce SLMs, NLMs, PLMs, and LLMs. Figure 1 provides a visual map of the history of the development of the LMs.
@@
  • Statistical Language Models: SLMs [67, 58, 68] originated in the 1990s as mathematical models addressing contextually relevant properties of natural language from a probabilistic statistical perspective. The essence of statistical language modeling lies in ascertaining the probability of a sentence occurring within a text. Considering $S$ as the sentence “I am very happy”, $P(\omega_{i})$ signifies the probability of the $i$-th word in the sentence: $\omega_{1}$ as “I”, $\omega_{2}$ as “am”, $\omega_{3}$ as “very”, and $\omega_{4}$ as “happy”. Now, the objective is to ascertain the likelihood of $S$ appearing in the text, denoted as $P(S)$:
@@
  • Neural Language Models: NLMs [69, 57, 56] leverage neural networks to predict the probabilities of subsequent words within sequences. They effectively handle longer sequences and mitigate the limitations associated with small $n$ in SLMs. Before delving into neural networks, let’s grasp the concept of word vectors. Humans effortlessly comprehend word meanings. For instance, “cat” and “dog” share closer semantic connections than “cat” and “tree” since they both represent animals. But how does a computer accomplish this? Computers operate using binary code — 0s and 1s — so human language needs translation into binary. Word vectors can accomplish this, which are numerical representations of human language, where each word corresponds to a distinct vector. These vectors usually possess fixed dimensions and can simulate word relationships through the angles between them. An intriguing example is the angle between the word vectors for “cat” and “dog” which is smaller than the angle between the word vector for “cat” and “tree”. Another illustrative example is “$Messi-Argentina+Portugal=Cristiano\ Ronaldo$”. Word2Vec [70] is a widely recognized tool for computing word vectors. Its function involves converting words into dense numerical representations, making words with similar semantics closer together in the vector space.
@@
    The efficacy of Word2Vec is notably influenced by its utilization of neural networks, which mirrors the structure and functionality of the human brain, comprising interconnected simple units known as neurons organized into different layers. Figure 2 provides a visualization of a simple NLM structure, composed of input, hidden, and output layers. Within this structure, each word vector $x_{t}\in\mathbb{R}^{m}$, the matrix $W_{xh}\in\mathbb{R}^{m(n-1)\times h}$, and the matrix $W_{ho}\in\mathbb{R}^{h\times|V|}$, where $h$ signifies the number of neurons in the hidden layer, and $V$ is the vocabulary containing all words recognized by the model. NLMs operate akin to the n-gram concept, assuming that the probability of each word depends only on its previous n-1 words. The first layer—the input layer—concatenates the word vectors of n-1 words, forming a vector $X\in\mathbb{R}^{m\times(n-1)}$. Subsequently, the second layer—the hidden layer—derives the output $H$ by applying a non-linear activation function, like sigmoid or tanh, to the matrix product of $X$ and $W_{xh}$. Following this, the third layer—the output layer—aims to forecast the subsequent word based on the hidden layer’s output. With $|V|$ neurons within the output layer, the resulting output vector $O\in\mathbb{R}^{|V|}$ is computed by multiplying $H$ and $W_{ho}$. This resultant vector then undergoes the Softmax function, producing a vector $O\textquoteright$ ’ containing the probability value assigned to each word. The Softmax function, concerning the output of the $i$-th neuron, is defined as follows:
    $O_{i}^{\prime}=\operatorname{Softmax}\left(O\textquoteright_{i}\right)=\frac{\exp\left(O_{i}\right)}{\sum_{j=1}^{|V|}\exp\left(O_{j}\right)},$ (4)
@@
    where $O_{i}^{\prime}$ denotes the output value of the $i$-th node, and $|V|$ represents the count of output nodes, corresponding to the classification categories. Utilizing the Softmax function allows the transformation of output values from multiclass classification into a probability distribution, ensuring a sum of 1 within the range [0, 1].
@@
  • Pre-trained Language Models: PLMs undergo initial training using an extensive volume of unlabeled text, enabling them to grasp fundamental language structures such as vocabulary, syntax, semantics, and logic — a phase termed pre-training. Subsequently, this comprehensive LM can be applied to various NLP tasks like machine translation, text summarization, and question-answering systems. To optimize its performance, models need to be trained a second time on a smaller dataset customized for a specific downstream task — a phase known as fine-tuning. This is the “pre-training and fine-tuning” learning paradigm. We can use a visual example to understand the “pre-training and fine-tuning”, as follows: in martial arts novels, a person who wants to become a martial arts master needs to have a solid foundation of internal martial arts, which can be acquired by training on a large variety of techniques. Then, the person can learn a specific skill, such as a sword or a palm strike, and master it quickly and effectively by applying internal martial arts principles. A large number of studies on PLMs have been built on this paradigm, which introduces different architectures [71, 61] (e.g., GPT-2 [9] and Bert [71]).
@@
  • Large Language Models: LLMs are trained on massive text corpora with tens of billions (or more) of parameters, such as GPT-3 [10], GPT-4 [11], PaLM [60], and LLaMA [72]. The goal of LLMs is to enable machines to understand human commands and adherence to human values. Their hallmark lies in the consolidation of two stages: initial pre-training on a vast general-purpose corpus followed by alignment with human values, rather than transitioning to a different domain. LLMs exhibit remarkable adaptability compared to PLMs, transitioning from specialized to general-purpose models. The substantial increase in model size, dataset volume, and computational prowess has resulted in significant enhancements across various tasks and unveiled remarkable capabilities absent in smaller models. For example, GPT-3 has the capability to leverage contextual information, a functionality that GPT-2 [9] lacks. This means that when GPT-3 is prompted with task-related examples, it utilizes them to enhance its problem-solving capabilities. The number of parameters for LLMs typically exceeds a hundred billion, and the training data is usually in the range of a few hundred GB to a few TB. A concrete example is that there are multiple versions of the GPT-2 model, with the largest version having 1.5 billion parameters and using 40GB of text data for training. In contrast, the largest version of the GPT-3 model has 175 billion parameters and uses 570GB of text data for training. This example illustrates the significant discrepancy between LLMs and PLMs concerning parameter count and training data volume.
@@ 2.2 Factors Propelling Large Language Models
  • Data Diversity: Data diversity has emerged as a crucial catalyst for the development of LLMs. Recent years have witnessed a surge in the availability and diversity of extensive internet-based data sources, furnishing these models with training data containing rich linguistic and worldly insights. The quality and origins of this data significantly influence the efficacy and capacity of LLMs. Earlier studies trained language models on a single domain of text, such as news articles [73], Wikipedia [30], or fiction books [74]. However, the proliferation of vast and heterogeneous text corpora across the web—including blogs, social media, forums, and reviews, among others—has spurred researchers to pivot towards training models on multi-domain texts. This shift enhances the model’s aptitude for generalization and multitasking. An exemplary instance is Meta AI’s open-source large language model, LLaMA [72], trained on publicly accessible datasets like CommonCrawl [^2] https://commoncrawl.org/get-started, C4 [62], Github [^3] https://github.com/, Wikipedia [^4] https://huggingface.co/datasets/wikipedia, Books [75], ArXiv [^5] https://arxiv.org/, and StackExchange [^6] https://stackexchange.com/.
@@
  • Computational Advancement: LLMs are neural network models with a large number of parameters. With the progression of deep learning (DL) technology in recent years, the scale and effectiveness of LLMs have notably escalated. Simultaneously, they present serious challenges due to their substantial computational demands [76]. Computational power signifies a computer system’s capability to execute computational tasks, often quantified in terms of floating-point operations per second. According to OpenAI calculations [^7] https://openai.com/research/ai-and-compute, since 2012, the amount of computation used for global AI training has shown exponential growth, doubling every 3.43 months on average, and now the amount of computation has expanded 300,000 times, far outpacing the rate of computational power growth. Fortunately, hardware innovators like Nvidia continually invent specialized devices such as GPUs, TPUs, and NPUs with amplified computational prowess, LLMs can be trained quickly. For instance, the launch of GPT-3 by OpenAI in May 2020 [10] underscores this reliance on robust hardware. Training GPT-3 on a single NVIDIA Tesla V100 GPU would require an estimated 355 years. Conversely, leveraging 1024×A100 GPUs, researchers estimated the potential to train GPT-3 in as little as 34 days. This exemplifies the indispensability of potent hardware in facilitating the development and training of LLMs.
@@
  • Algorithmic Innovation: Algorithmic innovation stands as the pivotal driving force behind LLMs, shaping their structure and functionality. From the initial rule- and statistics-based approach to the later DL-based approach, the algorithms of LMs have continuously evolved and improved. Presently, all LLMs are based on the transformer architecture [77], which employs a self-attention mechanism (see discussion in Section 3). This architecture offers distinct advantages over traditional recurrent [57] and convolutional neural networks [78], enabling fully parallel computation and adeptly capturing long-distance dependencies. Moreover, there exist numerous iterations and enhancements of the transformer, such as Transformer-XL [79], XLNet [80], ALBERT [81], each targeting specific aspects for optimization. These variations aim to enhance the attention mechanism, refine pre-training objectives, and curtail parameter counts, thereby advancing the transformer architecture in diverse dimensions.
3. PRINCIPLES AND TAXONOMY OF LARGE LANGUAGE MODELS
This section strives to clarify the fundamental principles and taxonomy of LLMs, intending to provide general practitioners with a more profound understanding of the operational mechanisms employed by these models and help readers choose the most suitable LLMs for their specific applications.